Determine the general solution to the lagrange equation

y2zp-x2zq=x2y


1
Expert's answer
2021-10-28T11:05:25-0400

"\displaystyle
y^2zp - x^2zq=x^2y\\
\text{Divide through by z to obtain}\\
y^2p-x^2q=\frac{x^2y}{z}\\
\text{The subsidiary equation is }\\
\frac{dx}{y^2} =\frac{dy}{-x^2}=\frac{zdz}{x^2y}\\
\text{Taking the the first equation}\\
\frac{dx}{y^2} =\frac{dy}{-x^2}\\
\text{Integrating both sides,we have}\\
\frac{x}{y^2}= -\frac{y}{x^2}+c_1\\
\implies c_1 = \frac{x}{y^2}+ \frac{y}{x^2}\\
\text{Next, we take another pair of ratios}\\
\frac{dx}{y^2}=\frac{zdz}{x^2y}\\
\text{Integrating, we have}\\
\frac{x}{y^2}=\frac{z^2}{2x^2y}+c_2\\
\implies c_2 = \frac{x}{y^2}-\frac{z^2}{2x^2y}\\
\text{Hence the general solution is}\\
\phi (\frac{x}{y^2}+ \frac{y}{x^2}), \frac{x}{y^2}-\frac{z^2}{2x^2y}=0"